# Rampa: Robotic Augmented Reality for Machine Programming and Automation

### v1

@@ V Conclusion
We introduce an AR-based framework, Rampa, that provides a comprehensive and intuitive platform for demonstrating robotic skill trajectories, integrating trajectory visualizations and modifications, real-time hand mimicry, and training and testing ML models. These functionalities of Rampa simplify the demonstration collection, an integral part of the PfD procedures, and enhance the human-robot interaction by leveraging AR. Importantly, we present quantitative evaluations of Rampa, supporting our initial hypothesis of increased effectiveness without sacrificing user experience or task precision. Thus, Rampa demonstrates the potential of leading to more accessible and efficient robotic programming for both novices and experts.
Currently, Rampa allows registering virtual objects while demonstrating trajectories. However, this study does not provide any evaluations of this functionality. The integration of this capability offers a more interactive and rich environment where users can use real and/or virtual objects to create their own scenarios for collecting demonstrations. Also, visualization of model-specific parameters - means and variances in the case of ProMPs - provides an enhanced introspection tool that can be used to ensure the robot’s behavior aligns with the desired outcome. Even though we use the UR10 robotic platform and ProMPs for convenience, we designed Rampa to be platform and model-agnostic. Therefore, the system can be (1) adapted to other robots as long as their URDF files are provided and (2) extended by integrating other ML models.
VI Acknowledgement
We would like to thank Emre Batuhan Goc and Muhammet Batuhan Ilhan for their contributions to the development of the underlying software of the Rampa system.

### v2

@@ VI Conclusion
We introduced an AR-based framework that provides a comprehensive and intuitive platform for demonstrating robotic skill trajectories, integrating trajectory visualizations and modifications, real-time hand mimicry, and training and testing ML models. These functionalities simplified the demonstration collection in PbD procedures and enhanced the human-robot interaction by leveraging AR. We presented quantitative evaluations, supporting our initial hypothesis of increased effectiveness without sacrificing user experience or task precision. Rampa illustrates the potential to lead to more accessible and efficient robotic programming for both novices and experts.